# Calibration-free single-frame super-resolution fluorescence microscopy

### v2

@@ Abstract
Molecular fluorescence microscopy is a leading approach to super-resolution and nanoscale imaging in life and material sciences. However, super-resolution fluorescence microscopy is often bottlenecked by system-specific calibrations and long acquisitions of sparsely blinking molecules. We present a deep-learning approach that reconstructs super-resolved images directly from a single diffraction-limited camera frame. The model is trained exclusively on synthetic data encompassing a wide range of optical and sample parameters, enabling robust generalization across microscopes and experimental conditions. Applied to dense terrylene samples with 150 ms acquisition time, our method significantly reduces reconstruction error compared to Richardson-Lucy deconvolution, ThunderSTORM multi-emitter fitting, and DECODE based on deep learning. The results confirm the ability to resolve emitters separated by 35 nm at 580 nm wavelength, corresponding to sevenfold resolution improvement beyond the Rayleigh criterion. By delivering unprecedented details from a single short camera exposure without prior information and calibration, our approach enables plug-and-play super-resolution imaging of fast, dense, or light-sensitive samples on standard wide-field setups.
@@ II Results and Discussion
We applied our calibration-free convolutional neural network (CFCNN) to fluorescence intensity images of terrylene molecules at various concentrations embedded in a polyvinyl alcohol (PVA) layer. These images were acquired using a fluorescence microscope, schematically shown in Fig. 1. The sample is illuminated by an excitation laser through a high numerical aperture oil-immersion objective lens, and the emitted fluorescence is collected by the objective and detected by an sCMOS camera.
@@
Due to the high molecule density and overlap of their PSFs, obtaining a super-resolved counterpart for the single dense camera image poses a significant challenge for traditional super-resolution methods. For details on the experiment, please refer to subsection A of the Methods.
Following the experimental setup, the camera image is processed by our CFCNN model. This deep-learning approach is calibration-free due to its unique methodology. The input to our model is a single intensity image, and the output is an upsampled, high-resolution reconstruction. The model is based on a convolutional neural network trained exclusively on synthetic data. This allows for the model to learn across a vast range of optical system parameters, covering configurations that would be impossible to obtain experimentally at such a scale, and eliminating systematic errors inherent in experimental measurements. Additionally, our network is not constrained to a single specific optical configuration. Instead, it is trained on diverse combinations of noise levels, PSF profiles, and other system parameters. This broad training scope enables the model to generalize effectively across various experimental conditions, making it robust both to variations in optical setup parameters across different systems and to variations within a single field of view, including aberrations, non-uniform background distributions, and other spatially varying changes. For further details on the CFCNN architecture and its training, please refer to the Methods, subsection B.
@@
Image 1 MAE $10^{-5}$] 4.64(8) 3.99* 3.0(2) 2.4(5)
KLD 6.0(2) 12.6* 13(2) 1(2)
Image 2 MAE $10^{-5}$] 3.98(7) 3.6(1) 3.0(1) 2.27(8)
KLD 7.8(7) 17.0(2) 16.3(5) 2.1(4)
Image 3 MAE $10^{-5}$] 3.52(3) 4.01(2) 3.99(4) 3.34(4)
KLD 10.0(3) 19.0(4) 25.9(2) 5.2(2)
@@
To validate and benchmark the performance and versatility of our CFCNN, we also employed three commonly used state-of-the-art methods on the same input images, specifically, the Richardson-Lucy (R-L) deconvolution algorithm as a reconstruction technique [37, 38], multi-emitter fitting (MEF) with ThunderSTORM (an ImageJ plugin) as a localization method [25], and DECODE as a deep learning approach [35]. All approaches were provided exclusively with a single resolution-limited image. The amount of additional information and parameter settings required by each method differs, as well as the necessary post-processing steps to produce the final super-resolved image. R-L algorithm iteratively applies Bayes theorem to update an estimate of the true image by maximizing the likelihood of the observed image. The algorithm can achieve significant resolution enhancement. However, it relies on accurate prior knowledge of the PSF profile, which necessitates either acquiring a calibration image of a sparse sample or the presence of an isolated molecule within the field of view. Besides being a time-consuming extra step, this also introduces a simplification, as the PSF shape may vary across the field of view. ThunderSTORM offers extensive customization options tailored to the user’s specific application. However, this also means the results are highly dependent on a large number of input parameters, such as the camera settings, image filtering, fitting algorithm method and parameters, etc. Their values need to be optimized for each given setup and measurement. The optimization involves balancing multiple trade-offs, requiring user expertise and adding complexity to the process. The output of the method is a table containing the emitter coordinates and intensities. Subsequent image rendering may lead to ambiguity since there are multiple ways to perform this step, and requires additional decisions in the processing pipeline. DECODE requires training on specific detailed calibration data, the acquisition of which is an experimentally demanding and time-consuming task. For the training, a fitted PSF $z$-stack needs to be provided, and parameters of the experiment need to be specified, such as the camera parameters, emitter density, brightness, etc. Once trained, it can be used only on data acquired under the same conditions, otherwise, retraining is necessary. Details on the settings for each method are given in part C of the Methods section. In contrast, our deep-learning approach operates in a plug-and-play manner, requiring no additional training or information beyond the input image, neither before nor after applying the CFCNN. The resulting super-resolved image outperforms the other methods across all evaluated metrics.
@@
The visual comparison between the input intensity image of fluorescent terrylene molecules, the reconstructed outputs from Richardson-Lucy deconvolution, multi-emitter fitting using ThunderSTORM, DECODE, and our CFCNN, is presented in Fig. 2. The output images are accompanied by the ground truth, obtained by leveraging the limited photostability of the molecules in PVA and their gradual localization from a recorded time-lapse of bleaching events (see the Methods section, part D, for details). We show the results for three independent experimental intensity images with different molecule concentrations and levels of background noise. The input image size is $50\times 50$ pixels and all outputs are upscaled to $200\times 200$ pixels to enhance the resolution and detail of the reconstructed image. The multi-emitter fitting, DECODE, and our manual localization technique for obtaining the ground truth all produce a list of localized emitters and their intensities, which are then rendered into the finer grid. The situation is different for R-L deconvolution, which directly returns an image of the original size. In this case, the input image was first upscaled via bicubic interpolation before applying the reconstruction algorithm. Finally, the CFCNN has upsampling layers incorporated in its architecture. Since the reconstructions consist of $200\times 200$ pixels containing small structures, they are convolved with a narrow Gaussian filter (with $\sigma=2$ px) for visualisation purposes in the figure. For further details on the methods employed and the process used to obtain the ground truth, please refer to the Methods section.
@@
Figure 3: Synthetic data with emitters arranged in the shape of a star. Images in panels (b)–(f) are convolved with a narrow Gaussian filter for better visualisation. (a) An example of the input image with $\textrm{SNR}=15.8$ obtained as the ground truth convolved with a Gaussian PSF with the width of 22 px on the $50\times 50$ pixel grid. (b) The image reconstruction by the Richardson-Lucy deconvolution algorithm. (c) The output of multi-emitter fitting using ThunderSTORM. (d) The resulting image produced by DECODE. (e) The reconstructed image by the CFCNN. (f) The real generated emitter distribution. (g) Mean absolute error between the outputs and the ground truth as a function of SNR for R-L (orange), MEF (green), DECODE (red), and the CFCNN (blue). (h) Kullback–Leibler divergence of the outputs from the ground truth as a function of SNR.
@@
To further investigate the methods performance beyond the experimental results, we conducted additional tests on synthetic datasets with controlled parameters and precisely known ground truth, which allows for accurate and reliable evaluation. Additionally, it enables a systematic study of the performance dependence on the intensity noise level. Throughout this work, we define the SNR as the difference between the maximum intensity and the average noise level, divided by the standard deviation of background noise. In Fig. 3 (a)–(e), we show a particular example of emitters arranged in the shape of a star. Panel (a) represents the input for the tested methods with the $\textrm{SNR}=15.8$, while panel (e) visualizes the generated ground truth with applied Gaussian filter. For the outputs of each studied method, also convolved with the narrow Gaussian filter, see panels (b)–(d). It is immediately apparent that the CFCNN reconstruction is virtually perfect, with all the emitters correctly positioned and their intensities accurately estimated. On the contrary, the R-L reconstruction contains numerous artificial emitters, not only in the vicinity of the true emitter positions but across the entire field of view. This makes it misleading and significantly complicates distinguishing genuine emitters from artifacts. The multi-emitter fitting struggles in dense regions and inaccurately estimates the emitter intensities, resulting in a compromised image. Some of the emitters were not resolved at all, leading to a loss of valuable information. Despite DECODE outperforming both R-L and MEF, it still fails to resolve several emitters in dense regions from a single camera image. We also evaluated MAE and KLD between the outputs and the well-defined ground truth. Fig. 3 (f) and (g) depict the dependence of these metrics on the noise level expressed by SNR for all three methods. Our CFCNN consistently outperforms the R-L algorithm, the MEF method, and DECODE across all SNR values.
@@
Futhermore, we generated a standalone dataset consisting of a pair of emitters with progressively increasing separation to assess the level of super-resolution achievable by our CFCNN. The followed procedure is outlined in the Methods section, part E. For the minimum resolvable separation between two emitters reconstructed by the CFCNN as a function of SNR, see Fig. 4. This separation $\delta_{R}$ is expressed in terms of the Rayleigh distance. It is evident that a significant level of super-resolution is attained even at the lowest SNR values. For higher SNR values, the achieved spatial resolution saturates towards $\delta_{R}=0.138$, corresponding to more than sevenfold enhancement. This saturation limit is imposed by the finite pixel size, and we anticipate that finer image sampling would further enhance the resolution. For the experimental images of terrylene molecules with an emission peak at $580$ nm, taken with a microscope objective with numerical aperture $1.4$, the CFCNN reaches the resolution of 35 nm. Furthermore, while SMLM typically achieves up to tenfold resolution improvement, it requires a long sequence of sparse images and meticulous calibration. By applying the CFCNN to each frame, we could reach the same level of super-resolution while significantly reducing the number of images needed, making the process considerably faster.
@@
So far, we have applied the CFCNN on samples with various SNR values and concentrations of molecules obtained from a high-NA microscope with a near-Gaussian PSF shape. To investigate the independence of the CFCNN on the PSF shape, we performed an additional analysis of low-NA microscopes. The corresponding PSFs were calculated using full vectorial electromagnetic simulations. In Fig. 5 (a) and (c) we show that the performance of our model is essentially independent of the objective numerical aperture, handling the transition from an Airy-type to a Gaussian PSF with the same level of accuracy. Together with the experimental results, this analysis clearly demonstrates that our CFCNN maintains its performance across a broad range of microscope configurations and PSF shapes.
Since optical setups are often affected by aberrations, we tested our CFCNN (trained only on ideal Gaussian and Airy PSFs) on aberrated PSFs to assess its robustness against such imperfections. In Fig. 5 (b) and (d), we present the image reconstruction quality in terms of MAE and KLD for increasing strengths $W$ of aberrations distorting the ideal PSF. Even though the CFCNN has never seen any aberrated PSF during training, it can handle a certain level of aberrations with only a slight deterioration in the metric values. Up to $0.15\lambda$, the MAE and KLD stay virtually constant for spherical aberration, defocus, and astigmatism, while increasing more for coma. These results highlight the generalization ability of the CFCNN to maintain high performance even under conditions for which it was never explicitly trained, demonstrating resilience to realistic optical imperfections.
@@ III Conclusion
Furthermore, we quantified the super-resolving ability of the CFCNN, achieving up to sevenfold enhancement relative to the standard Rayleigh limit. In our experiment with terrylene, with an emission wavelength of $580$ nm, this corresponds to resolving fluorophores separated by $35$ nm using a single camera frame with exposure time of 150 ms. The single-frame operation reduces measurement times down to tens of milliseconds, allowing real-time super-resolution imaging of dynamic and live samples that would be inaccessible by slower multi-frame techniques, and photosensitive specimens. Moreover, it relaxes the sparsity condition of blinking emitters, significantly broadening its applicability to a wider range of fluorescent markers. When combined with SMLM approaches, our CFCNN could notably shorten the overall acquisition time by substantially reducing the number of frames, while simultaneously improving accuracy. We also showed the CFCNN remains robust against moderate optical aberrations, despite being trained exclusively on ideal PSF shapes, highlighting the strong generalizability of the method.
The presented work demonstrates the potential of a calibration-free deep learning approach for rapid super-resolution fluorescence microscopy. Future improvements, such as incorporating additional upsampling layers for greater image magnification, introducing residual connections for enhancing the propagation of information [39], and using deeper network architectures, could further enhance the versatility and performance of our approach. With these advancements, we envision our network becoming a completely universal tool for high-precision microscopy, broadly applicable across numerous domains ranging from biomedical imaging to material sciences.

### v3

@@ Abstract
Molecular fluorescence microscopy is a leading approach to super-resolution and nanoscale imaging in life and material sciences. However, super-resolution fluorescence microscopy is often bottlenecked by system-specific calibrations and long acquisitions of sparsely blinking molecules. We present a deep-learning approach that reconstructs super-resolved images directly from a single diffraction-limited camera frame. The model is trained exclusively on synthetic data encompassing a wide range of optical and sample parameters, enabling robust generalization across microscopes and experimental conditions. Applied to dense terrylene samples with 150 ms acquisition time, our method significantly reduces reconstruction error compared to Richardson-Lucy deconvolution, ThunderSTORM multi-emitter fitting, and DECODE based on deep learning. The results confirm the ability to resolve emitters separated by 35 nm at 580 nm wavelength, corresponding to sevenfold resolution improvement beyond the Rayleigh criterion. Furthermore, we demonstrate strong generalization ability of the developed model and its resilience across a broad range of noise levels, numerical apertures, and optical aberrations. By delivering unprecedented details from a single short camera exposure without any prior information and calibration, our approach enables plug-and-play super-resolution imaging of fast, dense, or light-sensitive samples on common wide-field microscopy setups.
@@ II Results and Discussion
We applied our calibration-free convolutional neural network (CFCNN) to fluorescence intensity images of terrylene molecules at various concentrations embedded in a polyvinyl alcohol (PVA) layer. These images were acquired using a fluorescence microscope, schematically shown in Fig. 1. The sample is illuminated by an excitation laser through a high numerical aperture oil-immersion objective lens, and the emitted fluorescence is collected by the objective and detected by an sCMOS camera. For details on the experiment, please refer to subsection A of the Methods. Due to the high molecule density and overlap of their PSFs, obtaining a super-resolved counterpart for the single dense camera image poses a significant challenge for traditional super-resolution methods.
@@
Following the experimental setup, the camera image is processed by our CFCNN model. This deep-learning approach is calibration-free due to its unique methodology. The input to our model is a single intensity image, and the output is an upsampled, high-resolution reconstruction. The model is based on a convolutional neural network trained exclusively on synthetic data. This enables the model to learn across a vast range of optical system parameters, covering configurations that would be impossible to obtain experimentally at such a scale, and eliminating systematic errors inherent in experimental measurements. Additionally, our network is not constrained to a single specific optical configuration. Instead, it is trained on diverse combinations of noise levels, PSF profiles, and other system parameters. This broad training scope enables the model to generalize effectively across various experimental conditions, making it robust both to variations in optical setup parameters across different systems and to variations within a single field of view, including aberrations, non-uniform background distributions, and other spatially varying changes. For further details on the CFCNN architecture and its training, please refer to the Methods, subsection B, and our code and data public repository [37].
@@
Image 1 MAE $10^{-5}$] 4.64(8) 4.2(3) 3.6(5) 2.4(5)
KLD 6(2) 12(2) 14(5) 1(2)
Image 2 MAE $10^{-5}$] 3.98(7) 3.5(1) 4.49(7) 2.27(8)
KLD 7.8(7) 18(1) 22(1) 2.1(4)
Image 3 MAE $10^{-5}$] 3.52(3) 4.0(1) 4.58(6) 3.34(4)
KLD 10.0(3) 19.0(8) 31.2(8) 5.2(2)
@@
To validate and benchmark the performance and versatility of our CFCNN, we also employed three commonly used state-of-the-art methods on the same input images, specifically, the Richardson-Lucy (R-L) deconvolution algorithm as a reconstruction technique [38, 39], multi-emitter fitting (MEF) with ThunderSTORM (an ImageJ plugin) as a localization method [25], and DECODE as a deep learning approach [35]. All approaches were provided exclusively with a single resolution-limited image. The amount of additional information and parameter settings required by each method differs, as well as the necessary post-processing steps to produce the final super-resolved image. R-L algorithm iteratively applies Bayes theorem to update an estimate of the true image by maximizing the likelihood of the observed image. The algorithm can achieve significant resolution enhancement. However, it relies on accurate prior knowledge of the PSF profile, which necessitates either acquiring a calibration image of a sparse sample or the presence of an isolated molecule within the field of view. Besides being a time-consuming extra step, this also introduces a simplification, as the PSF shape may vary across the field of view. ThunderSTORM offers extensive customization options tailored to the user’s specific application. However, this also means the results are highly dependent on a large number of input parameters, such as the camera settings, image filtering, fitting algorithm method and parameters, etc. Their values need to be optimized for each given setup and measurement. The optimization involves balancing multiple trade-offs, requiring user expertise and adding complexity to the process. The output of the method is a table containing the emitter coordinates and intensities. Subsequent image rendering may lead to ambiguity since there are multiple ways to perform this step, and requires additional decisions in the processing pipeline. DECODE requires training on specific detailed calibration data, the acquisition of which is an experimentally demanding and time-consuming task. For the training, a fitted PSF $z$-stack needs to be provided, and parameters of the experiment must be specified, such as the camera parameters, emitter density, brightness, etc. Once trained, it can be used only on data acquired under the same conditions, otherwise, retraining is necessary. Details on the settings for each method are given in part C of the Methods section. In contrast, our deep-learning approach operates in a plug-and-play manner, requiring no additional training or information beyond the input image, neither before nor after applying the CFCNN. The resulting super-resolved image outperforms the other methods across all evaluated metrics.
@@
The visual comparison between the input intensity image of fluorescent terrylene molecules, the reconstructed outputs from Richardson-Lucy deconvolution, multi-emitter fitting using ThunderSTORM, DECODE, and our CFCNN, is presented in Fig. 2. The output images are accompanied by the ground truth, obtained by leveraging the limited photostability of the molecules in PVA and their gradual localization from a recorded time-lapse of bleaching events (see the Methods section, part D, for details). We show the results for three independent experimental intensity images with different molecule concentrations and levels of background noise. The input image size is $50\times 50$ pixels and all outputs are upscaled to $200\times 200$ pixels to enhance the resolution and detail of the reconstructed image. The multi-emitter fitting, DECODE, and our manual localization technique for obtaining the ground truth all produce a list of localized emitters and their intensities, which are then rendered into the finer grid. The situation is different for R-L deconvolution, which directly returns an image of the original size. In this case, the input image was first upscaled via bicubic interpolation before applying the reconstruction algorithm. Finally, the CFCNN has upsampling layers incorporated in its architecture. Since the reconstructions consist of $200\times 200$ pixels containing small structures, they are convolved with a narrow Gaussian filter (with $\sigma=2$ px) for visualization purposes in the figure. For further details on the methods employed and the process used to obtain the ground truth, please refer to the Methods section.
@@
Figure 3: Synthetic data with emitters arranged in the shape of a star. Images in panels (b)–(f) and (h)–(l) are convolved with a narrow Gaussian filter for better visualization. (a, g) Examples of the input image with $\textrm{SNR}=15.8$ and $\textrm{SNR}=125.9$, respectively, obtained as the ground truth (f, l) convolved with a Gaussian PSF with the width of 22 px on the $50\times 50$ pixel grid and the corresponding noise backgrounds added. (b, h) The image reconstructions by the Richardson-Lucy deconvolution algorithm. (c, i) The outputs of multi-emitter fitting using ThunderSTORM. (d, j) The resulting images produced by DECODE. A uniquely trained model was used for each SNR. (e, k) The reconstructed images by the CFCNN without recalibration and retraining. (m) Mean absolute error between the outputs and the ground truth as a function of SNR for R-L (orange), MEF (green), DECODE (red), and the CFCNN (blue). (n) Kullback–Leibler divergence of the outputs from the ground truth as a function of SNR. In the case of DECODE, a separate model was trained on the corresponding SNR and PSF for each data point.
@@
To further investigate the methods performance beyond the experimental results, we conducted additional tests on synthetic datasets with controlled parameters and precisely known ground truth, which allows for accurate and reliable evaluation. Additionally, it enables a systematic study of the performance dependence on the intensity noise level. Throughout this work, we define the SNR as the difference between the maximum intensity and the average noise level, divided by the standard deviation of background noise. In Fig. 3 (a)–(l), we show a particular example of emitters arranged in the shape of a star. Panel (a) represents the input for the tested methods with the $\textrm{SNR}=15.8$, while panel (f) shows the generated ground truth with the applied narrow Gaussian filter for better visualization. For the outputs of each studied method, also convolved with the narrow Gaussian filter, see panels (b)–(e). The second row showcases the input image with $\textrm{SNR}=125.9$ in panel (g) and the corresponding results in (h)–(k). It is immediately apparent that the CFCNN reconstruction is virtually perfect, with all the emitters correctly positioned and their intensities accurately estimated, except for a slightly weaker rendering of the central emitter for $\textrm{SNR}=15.8$. On the contrary, the R-L reconstruction contains numerous artificial emitters, not only in the vicinity of the true emitter positions but across the entire field of view, particularly for the low SNR regime. The same also holds for the high SNR, but the artifacts are less bright. This renders the R-L results misleading and significantly complicates distinguishing genuine emitters from artifacts. The multi-emitter fitting struggles in dense regions and inaccurately estimates the emitter intensities, resulting in a compromised image. Some of the emitters were not resolved at all, leading to a loss of valuable information. Finally, DECODE provides a reasonable output for $\textrm{SNR}=15.8$, however, with several emitters slightly misplaced and their intensities imprecisely estimated. For $\textrm{SNR}=125.9$ and other SNR values in the range approximately from 30 to 300, DECODE omits the entire central part of the star sample despite being trained for the specific SNR and PSF values. The main reason for this failure is the extremely low generalization ability of its trained models. As the density of emitters, their brightness, and SNR must be specified during the training stage, the resulting model works well only for images that match exactly those parameters. Any mismatch, even local, in a part of the field of view, results in significant errors in the DECODE output. These errors typically manifest as missing emitters, likely due to the relatively strong sparsity enforcement in DECODE training. Generally speaking, MEF and DECODE tend to omit emitters, which may represent tolerable errors in single-molecule localization microscopy, such as STORM and PALM, where the density of active molecules in a single frame is low and there is a high probability of repeated molecule localization in several frames. However, when super-resolution is performed from a single camera image, these methods fail to provide high-fidelity super-resolution imaging, especially for images with inhomogeneous emitter concentrations across the field of view.
We also evaluated MAE and KLD between the outputs and the well-defined ground truth to numerically quantify the reconstruction performance. Fig. 3 (m) and (n) depict the dependence of these metrics on the noise level expressed by SNR for all four methods. As before, R-L, MEF, and DECODE were supplied with all necessary information. Specifically, for DECODE, each data point corresponds to a unique model trained for the respective SNR value, requiring tedious recalibration and retraining that takes dozens of hours of extra work. Our CFCNN consistently outperforms the R-L algorithm, the MEF method, and DECODE across all SNR values, despite operating without any additional information, calibration, or retraining.
@@
Furthermore, we generated a standalone dataset consisting of a pair of emitters with progressively increasing separation to assess the level of super-resolution achievable by our CFCNN. The followed procedure is outlined in the Methods section, part E. For the minimum resolvable separation between two emitters reconstructed by the CFCNN as a function of SNR, see Fig. 4. This separation $\delta_{R}$ is expressed in terms of the Rayleigh distance. It is evident that a significant level of super-resolution is attained even at the lowest SNR values. For higher SNR values, the achieved spatial resolution saturates towards $\delta_{R}=0.138$, corresponding to more than sevenfold enhancement. This saturation limit is imposed by the finite pixel size, and we anticipate that finer image sampling would further enhance the resolution. For the experimental images of terrylene molecules with an emission peak at $580$ nm, taken with a microscope objective with numerical aperture $1.4$, the CFCNN reaches the resolution of 35 nm. Furthermore, while SMLM typically achieves up to tenfold resolution improvement, it requires a long sequence of sparse images and meticulous calibration. By applying the CFCNN to each frame, we could reach the same level of super-resolution while significantly reducing the number of images needed, making the process considerably faster.
@@
So far, we have applied the CFCNN on samples with various SNR values and concentrations of molecules obtained from a high-NA microscope with a near-Gaussian PSF shape. To investigate the independence of the CFCNN on the PSF shape, we performed an additional analysis of low-NA microscopes. The corresponding PSFs were calculated using full vectorial electromagnetic simulations. In Fig. 5 (a) and (c), we show that the performance of our model is essentially independent of the objective numerical aperture, handling the transition from an Airy-type to a Gaussian PSF with the same level of accuracy. This is particularly striking because the CFCNN model was not trained on these intermediate NA values. Together with the experimental results, this analysis clearly demonstrates that our CFCNN maintains its performance across a broad range of microscope configurations and PSF shapes, showing the unprecedented degree of generalization.
Since optical setups are often affected by aberrations, we tested our CFCNN (trained only on ideal Gaussian and Airy PSFs) on aberrated PSFs to assess its robustness against such imperfections. In Fig. 5 (b) and (d), we present the image reconstruction quality in terms of MAE and KLD for increasing strength $W$ of aberrations distorting the ideal PSF. Even though the CFCNN has never seen any aberrated PSF during training, it can handle a certain level of aberrations with only a slight deterioration in the metric values. Up to $0.15\lambda$, the MAE and KLD stay virtually constant for spherical aberration, defocus, and astigmatism, while increasing more for coma. These results again highlight the generalization ability of the CFCNN to maintain high performance even under conditions for which it was never explicitly trained, demonstrating resilience to realistic optical imperfections. This resilience can be increased even further by expanding the CFCNN training dataset to aberrated PSFs and, potentially, to other imperfections at the expense of increased computational cost. Finally, it is important to stress that the CFCNN training is performed only once, and the trained model is then readily applicable to any samples and microscope setups without retraining, which represents a completely novel approach to deep-learning super-resolution and significant simplification of any application workflow.
@@ III Conclusion
Furthermore, we quantified the super-resolving ability of the CFCNN, achieving up to sevenfold enhancement relative to the standard Rayleigh limit. In our experiment with terrylene, with an emission wavelength of $580$ nm, this corresponds to resolving fluorophores separated by $35$ nm using a single camera frame with an exposure time of 150 ms. The single-frame operation reduces measurement times to dozens of milliseconds, allowing real-time super-resolution imaging of dynamic and live samples that would be inaccessible with slower multi-frame techniques, and photosensitive specimens. Moreover, it relaxes the sparsity condition of blinking emitters, significantly broadening its applicability to a wider range of fluorescent markers. When combined with SMLM approaches, our CFCNN could notably shorten the overall acquisition time by substantially reducing the number of frames, while simultaneously improving accuracy. We also showed that the CFCNN remains robust against moderate optical aberrations, despite being trained exclusively on ideal PSF shapes, highlighting the strong generalizability of the method.
The presented work demonstrates the potential of a calibration-free deep learning approach for rapid super-resolution fluorescence microscopy. Future improvements, such as incorporating additional upsampling layers for even higher resolution, introducing residual connections for enhancing the propagation of information [40], and using deeper network architectures, could further enhance the versatility and performance of our approach. With these advancements, we envision our network becoming a completely universal tool for high-precision microscopy, broadly applicable across numerous domains ranging from biomedical imaging to material sciences.